A vessel at rest explodes, breaking into three pieces. Two pieces, having equal mass, fly off perpendicular to one another with the same speed of 30 m/s. The third piece has three times the mass of each other piece. What is the direction and magnitude of its velocity immediately after the explosion?

Grade:12th pass

1 Answers

Khimraj
5 years ago
As external force is zero.
apply conservation of momentum.
m*0 = (m/5)*30\hat{i} + (m/5)*30\hat{j} + (3m/5)*v
so v = -10\hat{i} -10\hat{j}
magnitude = 10\sqrt{2}
Hole it clears............................................